Show that the velocity of impact at the ground of an object thrown

vertically upward is equal to the initial velocity of the upward throw.


1
Expert's answer
2021-03-21T11:26:12-0400

From the law of conservation of energy the total energy of an object remains constant.

At the initial moment the total energy is equal to the kinetic energy:


"E_k=\frac{mv_0^2}{2},"


where m - the mass of object;

"v_0" - the initial velocity of the upward throw.


At the highest point of trajectory the kinetic energy is zero and the total energy is equal to the potential energy:


"E_p=mgh,"


where h - the altitude of object.


Then the object begins to drop downward and on graond level its total enrrgy is equal to he kinetic energy again:


"E_k=\frac{mv_i^2}{2},"


where "v_i" - the velocity of impact at the ground.


So, "\frac{mv_i^2}{2}=\frac{mv_0^2}{2},"


and "v_i=v_0."